# Diversifying the Expert Knowledge for Task-Agnostic Pruning in Sparse Mixture-of-Experts

### v3

@@ 4.1 Experiment Setup
Studied Models. We take pruning three MoE-based architectures as an example, including the Mixtral, Deepseek, and Qwen. For the Mixtral architecture, the Mixtral-8x7B has $32$ sparse MoE-involved layers, in each there are $8$ experts. The Mixtral-8x22B is similar to Mixtral-8x7B but with $56$ sparse MoE layers. During the inference, each token will select $2$ experts in each MoE layer. The deepseek model has $28$ layers, and there are $64$ experts in each layer. Each token will pass $2$ shared experts and select $6$ experts during the inference. The Qwen model also has $28$ MoE layers with $64$ experts in each layer but will activate $8$ experts during the inference.
@@ 4.3 Pruning the Deepseek Model
We also evaluate the effectiveness of our proposed pruning method on compressing the DeepSeek architecture. Unlike Mixtral-MoE, DeepSeek-MoE features a shared expert and incorporates more fine-grained experts at each MoE layer. Specifically, we pruned the non-shared experts of DeepSeek-MoE-16B, reducing the number of experts from 64 to 48 using various model pruning strategies. The results are illustrated in fig. 4***We only show the pruning results of the Deepseek and Qwen models on the most challenging MMLU task and the average performance across MMLU, BoolQ, OpenBookQA, and RTE. Full results can be found in supplementary..
@@ 4.4 Pruning the Qwen Model
The Qwen architecture is also utilized in our experiments. We study the Qwen2-57B-14A to evaluate the performance of different model pruning strategies. We prune the experts from $64$ experts to $48$ experts in each layer and evaluate the performance on different tasks. The results are reported in fig. 5.
When pruning $\frac{1}{3}$ experts of the full model, the data-centric strategy delivers the best performance, outperforming both our model-centric method and the count-guided baseline method. On the MMLU task, our model-centric strategy achieves approximately $1.5\%$ better performance compared to the count-guided pruning strategy.
4.5 Empirical Analysis on Expert Hints
While expert pruning brings a significant speedup by reducing the parameters, what changes are introduced to different MoE layers? To answer this question, we analyze the change of expert hints on the calibration dataset C4.
Figure 6: Statistics of the visiting frequency for all experts in different MoE layers.
The statistic results are shown in fig. 6. We can see that in most MoE layers in Mixtral-8x7B, many pairs of experts have similar hints. In contrast, pruning differentiates the hints of experts. Compared to directly using hints as the pruning goal, using expert knowledge as the pruning criterion results in more significant changes in hints. Additionally, the merging operation on the gate aggregates the new expert (last column) more tokens, further increasing the hint differences.
@@ 4.6 Ablation Study and Discussion
Comparison with the greedy search. To further demonstrate the superiority of our proposed expert pruning strategy, we employ a greedy search approach to identify the optimal candidates in each MoE layer for pruning and compare the results with those obtained using our method. Specifically, we enumerate all possible combinations of removing two experts in each MoE layer, evaluate the pruned models on the MMLU task, and record the best combination for each layer. We then merge these results to prune the model across all layers. The model pruned using the greedy search approach achieves a performance of $62.22\%$, while our data-centric strategy achieves $61.19\%$, showing a comparable result. We provide the full enumeration results in appendix C.
On the used samples. We conduct experiments to study the effect of both sample size and augmentation in data-centric pruning methods. As shown in table 3, increasing the number of samples from $128$ to $512$ without augmentation leads to a noticeable performance improvement of up to $0.6\%$. However, when the sample size is further increased to $2,048$, the average performance drops significantly to $61.6\%$. In contrast, when comparing this result with table 1, where $128$ samples were used with augmentation, we observe an average performance improvement of $0.3\%$. This demonstrates the importance of both sample size and augmentation in enhancing generalization during the pruning process, as discussed in section 3.4.
We attribute the performance drop with large sample sizes to the mean operation in our expert similarity estimation. While the mean preserves the feature dimension for efficient pairwise distance calculations, it reduces the precision of similarity estimates due to "round-off" errors, compared to per-sample metrics like CKA used in section 3.2.
@@
5 Conclusion and Future Work
@@
However, we didn’t account for the uniqueness of different MoE layers during the pruning process, often setting the same number of experts to be reduced across all layers. Our analysis reveals that some experts have low visitation frequencies, indicating a significant potential for further pruning.
@@ Appendix A Experiment details
We conducted our experiments using four 80GB NVIDIA A100 GPUs. In our learning strategy, we used SGD as the optimizer to learn the coefficients for expert merging, with the learning rate set to $1\times 10^{-3}$. The coefficients were initialized as identity matrices and optimized over $50$ epochs. For baseline methods and models, all usage and distribution comply with the terms of their license, i.e., Mixtral (Apache License 2.0), Deepseek (MIT license), Qwen (Tongyi Qianwen license), lm evaluation harness (MIT license).

### v4

@@ 4.1 Experiment Setup
Studied Models. We take pruning three MoE-based architectures as an example, including the Mixtral, Deepseek, and Qwen. For the Mixtral architecture, the Mixtral-8x7B has $32$ sparse MoE-involved layers, in each there are $8$ experts. The Mixtral-8x22B is similar to Mixtral-8x7B but with $56$ sparse MoE layers. During the inference, each token will select $2$ experts in each MoE layer. The deepseek model has $28$ layers, and there are $64$ experts in each layer. Each token will pass $2$ shared experts and select $6$ experts during the inference. The Qwen model also has $28$ MoE layers with $64$ experts in each layer but will activate $8$ experts during the inference. We include more experimental details in appendix A.
@@ 4.3 Pruning the Deepseek Model
We also evaluate the effectiveness of our proposed pruning method on compressing the DeepSeek architecture. Unlike Mixtral-MoE, DeepSeek-MoE features a shared expert and incorporates more fine-grained experts at each MoE layer. Specifically, we pruned the non-shared experts of DeepSeek-MoE-16B, reducing the number of experts from 64 to 48 using various model pruning strategies. The results are illustrated in fig. 4We only show the pruning results of the Deepseek and Qwen models on the most challenging MMLU task and the average performance across MMLU, BoolQ, OpenBookQA, and RTE. Full results can be found in supplementary..
@@ 4.4 Pruning the Qwen Model
The Qwen architecture is also utilized in our experiments. We study the Qwen2-57B-14A to evaluate the performance of different model pruning strategies. We prune the experts from $64$ experts to $48$ experts in each layer and evaluate the performance on different tasks.
Reported in fig. 5, when pruning $\frac{1}{3}$ experts of the full model, the data-centric strategy delivers the best performance, outperforming both our model-centric method and the count-guided baseline method. On the MMLU task, our model-centric strategy achieves approximately $1.5\%$ better performance compared to the runner-up one.
Performance discrepancy between Deepseek and Qwen.
The difference in performance appears to stem from variations in model architectures and training strategies. Specifically, Deepseek-MoE-16B uses shared experts to learn a broad range of knowledge, which can lead to highly diverse expert functions. Consequently, identifying similar experts for pruning becomes challenging, and removing them risks losing unique knowledge—resulting in notable performance drops. In contrast, Qwen2-57B-14A initializes its expert weights from a pre-trained dense model, increasing expert similarity post-training. Meanwhile, Deepseek-MoE-16B’s experts are randomly initialized and trained from scratch, resulting in fewer interchangeable experts. Pruning such independently specialized experts leads to a more pronounced performance decline.
@@ 4.5 Ablation Study and Discussion
Comparison with the greedy search. To further demonstrate the superiority of our proposed expert pruning strategy, we employ a greedy search approach to identify the optimal candidates in each MoE layer for pruning and compare the results with those obtained using our method. We enumerate all possible combinations of removing two experts in each layer, evaluate the pruned models on the MMLU task, and record the best combination for each layer. We then merge these results to prune the model across all layers. The model pruned using the greedy search approach achieves a performance of $62.22\%$, while our data-centric strategy achieves $61.19\%$, showing a comparable result. Full results are shown in appendix C.
On the used samples. We conduct experiments to study the effect of both sample size used for the calibration dataset and augmentation in data-centric pruning methods. As shown in table 3, increasing the number of samples from $128$ to $512$ without augmentation leads to a noticeable performance improvement of up to $0.6\%$.In contrast, when comparing this result with table 1, where $128$ samples were used with augmentation, we observe an average performance improvement of $0.3\%$. This demonstrates the importance of both sample size and augmentation in enhancing generalization during the pruning process, as discussed in section 3.5.
@@
Efficiency improvement after MoE pruning To demonstrate the benefits of MoE pruning, we provide a statistical efficiency analysis for Mixtral-8x7B and DeepSeek-MoE-16B. We set the batch size to 8 and evaluate the performance on the MMLU task using the lm_eval library (Gao et al., 2021). Other settings are consistent with the previous experiment. As shown in table 5, pruning two experts from each MoE layer in the Mixtral model achieves a 1.17× speedup and a 16.8% reduction in GPU memory usage. Pruning 16 experts from each MoE layer in the Deepseek model reduces inference time by 15.5% and GPU memory usage by 14.1%. These results highlight the computational benefits of expert pruning across different model scales. More analysis can be found in appendix F.
Table 5: Computation cost of evaluating Mixtral-8x7B and DeepSeek-MoE-16B on the MMLU task. Experiment settings are consistent with our paper.
Model # Experts Time (s) Mem (GB)
Mixtral-8x7B 8 (Ori.) 281 125
6 241 104
4 223 83
[…]
60 446 61
56 429 59
52 413 57
48 386 55
5 Conclusion
@@ Appendix A Experiment details
We conducted our experiments using four 80GB NVIDIA A100 GPUs.
In our learning strategy for expert merging, we used SGD as the optimizer to learn the coefficients for expert merging, with the learning rate set to $1\times 10^{-3}$. We randomly sample $128$ samples from the calibration set (C4) and partition them into a $3:1$ ratio as the training and evaluation sets, respectively. The coefficients were initialized as identity matrices and optimized over $50$ epochs until they converged.
For baseline methods and models, all usage and distribution comply with the terms of their license, i.e., Mixtral (Apache License 2.0), Deepseek (MIT license), Qwen (Tongyi Qianwen license), lm evaluation harness (MIT license). We use Copilot to help with debugging and coding.
@@ Appendix E Full results on pruning Deepseek and Qwen
Appendix F Empirical Analysis on Expert Hints
We analyze the change of expert hints on the calibration dataset C4, where the expert hint refers to the visiting frequency of the expert on the calibration set.
Figure F6: Statistics of the visiting frequency for all experts in different MoE layers.
The statistic results are shown in fig. F6. We can see that in most MoE layers in Mixtral-8x7B, many pairs of experts have similar hints. In contrast, pruning differentiates the hints of experts. Compared to directly using hints as the pruning goal, using expert knowledge as the pruning criterion results in more significant changes in hints. Additionally, the merging operation on the gate aggregates the new expert (last column) more tokens, further increasing the hint differences.
Appendix G Pseudo-code implementation
Algorithm 1 Pruning algorithm for given MoE layer
1:The set of experts =$\mathcal{F}=\{f_{1},f_{2},...,f_{N}\}$ in the given MoE layer to prune, each expert $f_{i}$ consisting of three layers parameterized by $\theta^{i}_{1},\theta^{i}_{2},\theta^{i}_{3}$, respectively, the number of experts to reserve $r$, the calibration dataset $\mathcal{D}$, strategy s.
2:Pruned MoE layer with reduced experts $\{f_{1},f_{2},...,f_{r}\}$.
3:function MoE pruning(MoE layer $\mathcal{F}$, number of experts $r$ to reserve)
4:     Initialize $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathcal{A}\}$.
5:     for each $i$ in $N$ do
6:         if s==Data-centric then
7:              $\mathcal{R}(f_{i})=f_{i}(\text{Mixup}(\mathcal{D}))$.
8:         else if s==Vectorized weights then
9:              $\mathcal{R}(f_{i})=\{\theta^{i}_{1},\theta^{i}_{2},\theta^{i}_{3}\}$
10:         else if s==Surrogate weights then
11:              $\mathcal{R}(f_{i})=\theta^{i}_{2}(\theta^{i}_{1}\cdot\theta^{i}_{3}$)
12:     for each $i$ in $N$ do
13:         for each $j$ in $N$ do
14:              $\mathcal{A}(\mathcal{E}_{ij})=\text{CKA}(\mathcal{R}(f_{i}),\mathcal{R}(f_{j}))$
15:     Optimize eq. 4 on $\mathcal{G}$ to find $r$ subgroup of experts.
16:     Merge experts with their routers clustered within the same subgroup based on eq. 5.
17:     return Pruned $\mathcal{F}$
Solving eq. 4 is an NP-hard problem. To obtain an approximate solution, two methods can be employed. One approach is to use spectral clustering (Ng et al., 2001), while the other is a greedy strategy that iteratively merges the pair of experts with the highest similarity to construct the pruned expert graph. Although both methods yield comparable performance, the greedy approach is significantly more efficient, particularly when dealing with large language models that have many experts per MoE layer. In practice, for data-centric method, we also observe that using cosine similarity performs on par with centered kernel alignment (CKA), while being more computationally efficient for measuring expert similarity. At the same time, the mean-squared error also works well, which we adopt in the model-centric method.
Appendix H Discussion on data- and model-centric methods
The data-centric method is a natural extension of the paper’s starting point. However, its implementation requires access to a small portion of the pre-training dataset or a surrogate dataset with a similar distribution, which may pose practical challenges.
On the one hand, achieving good distributional similarity between the surrogate and real datasets is indeed crucial for the success of the data-centric method. Without such similarity, the method can fail entirely. In our trials, we tried using random inputs to identify similar experts and prune the MoE model, which led to the crashed performance of the pruned experts.
On the other hand, while surrogate datasets are generally simpler than real-world datasets, they still capture critical characteristics learned by different experts. Notably, no pre-trained data is publicly available for the models we study, including Mixtral, DeepSeek, and Qwen. However, leveraging a small portion of the open-source pre-trained dataset C4 still yields comparable performance on diverse benchmarks, as shown in our results.
In contrast, the model-centric method identifies similar experts directly by analyzing the model’s weights, bypassing the need for any data. This approach builds on the observation that training on the same dataset produces similar weight distributions, which can serve as a surrogate for predicting outputs given the same inputs. In the context of MoEs, when the gating mechanism consistently routes similar tokens to a group of experts, those experts tend to exhibit similar weights. Therefore, the model-centric strategy aligns with the paper’s initial data-dependent premise by offering a complementary perspective on expert similarity without relying on external datasets.